Find the equation of the hyperbola with foci at (1,2) (1,6) and vertices at (1,3) (1,5)


1
Expert's answer
2021-09-20T14:54:31-0400

Foci "(h, k\pm c),"

Vertices "(h, k\pm a)"

Equation


"\dfrac{(y-k)^2}{a^2}-\dfrac{(x-h)^2}{b^2}=1"

"a^2+b^2=c^2"

Foci "(1,2),(1,6)"

Vertices "(1,3),(1,5)"

Substitute


"h=1, k-c=2, k+c=6"

"h=1, k-a=3, k+a=5"

Then "h=1, k=4, c=2, a=1"


"b^2=c^2-a^2=(2)^2-(1)^2=3"

The equation of the hyperbola is


"\dfrac{(y-4)^2}{1}-\dfrac{(x-1)^2}{3}=1"